I have a question about finding midpoints of a line segments when given the end points example: {-4/3,12/7},{7/4,2}
Answer by checkley77(12844)

{-4/3,12/7},{7/4,2}
(7/4+4/3)/2
[(3*7+4*4)/12]/2
[(21+16)/12]/2
37/12*1/2=37/24 ANS. FOR THE X COORDINATE.
[(2-12/7)/2]
[(14-12/7)/2]
2/7*1/2=2/14=1/7 ANS. FOR THE Y COORDINATE.
(37/24,1/7)